# On the geometry of spaces of filtrations on local rings

### v1

@@ 4.1. Supnorm metrics and homogenization
Throughout this section we will use the language of norms.
@@
Fix a norm $\rho\in\mathcal{N}^{h}$. We define a function $d_{\infty,\rho}:\mathcal{N}\times\mathcal{N}\to\mathbb{R}_{\geq 0}$ by
@@
By definition, $d_{\infty,\rho}$ is symmetric and non-negative for any $\rho$. Given $\chi,\chi^{\prime},\chi^{\prime\prime}\in\mathcal{N}$, since $|\chi(f)-\chi^{\prime\prime}(f)|\leq|\chi(f)-\chi^{\prime}(f)|+|\chi^{\prime}(f)+\chi^{\prime\prime}(f)|$, the triangle inequality holds. This proves that $d_{\infty,\rho}$ is a pseudometric.
The second claim follows from the fact that given $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\in\mathbb{R}_{>0}$ such that $M^{-1}\rho\leq\rho^{\prime}\leq M\rho$. ∎
@@
Denote $d:=d_{\infty}(\widehat{\chi},\widehat{\chi}^{\prime})$. By Lemma 4.3, we may assume that $d=\sup\frac{\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)}{\chi_{0}(f)}$. For any $\epsilon>0$, there exits $f\in\mathfrak{m}$ such that
@@
4.2. Locally Lipschitz continuity of log canonical thresholds
In this section we prove that the log canonical thresholds is locally Lipschitz with respect to the $d_{\infty}$-topology.
@@ 4.2. Locally Lipschitz continuity of log canonical thresholds
Since $\mathrm{lct}(\chi)=\mathrm{lct}(\widehat{\chi})$ for any $\chi\in\mathcal{N}$, by Lemma 4.4, we may assume that $\chi,\chi^{\prime}\in\mathcal{N}^{h}$. Fix $c=c(\chi):=\inf_{f\in\mathfrak{m}}\frac{\chi(f)}{\chi_{0}(f)}>0$ and denote $a:=\mathrm{lct}(\chi)>0$.
@@
$\frac{v(f)}{\chi^{\prime}(f)}=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi(f)-\chi^{\prime}(f)}{\chi^{\prime}(f)}\right)$ (4.1)
@@ 4.3. Comparison of different topologies on $\mathrm{Fil}^{s}$
Now we would like to compare the topologies on the space $\mathrm{Fil}^{s}$.  [^4] Since the product topology is in general not first countable and hence not obviously sequential, we avoid using the term finer for them, but only compare convergence for sequences. First we show that the a $d_{1}$-converging seqeunce converges in the $+$-topology.
@@
Let $\mathfrak{a}_{\bullet}$,$\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ for $k\in\mathbb{Z}_{>0}$. If $\lim_{k}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet})=0$, then $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}$ in the $+$-topology.
@@
We claim that if $\mathfrak{a}_{\bullet,k}\subset\mathfrak{b}_{\bullet,k}$ satisfy $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{b}_{\bullet,k})=0$, then for any $v\in\mathrm{Val}^{+}_{X,x}$, we have $\lim_{k\to\infty}(v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{b}_{\bullet,k}))=0$.
@@
Now it remains to prove the claim, which follows essentially from the same argument as [BLQ24, Proosition 3.12]. We include a sketch for the reader’s convenience. Assume to the contrary that there exists $\epsilon>0$ and $v\in\mathrm{Val}^{+}_{X,x}$ such that $v(\mathfrak{a}_{\bullet,k})>v(\mathfrak{b}_{\bullet,k})+\epsilon$ for any $k$. Then for $k\in\mathbb{Z}_{>0}$, we may choose $\ell_{k}\in\mathbb{Z}_{>0}$ and $f_{k}\in\mathfrak{b}_{\ell_{k},k}$ such that
@@
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\leq M\cdot d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}).$ (4.2)
@@
To summarize, we have the following diagram for convergence of sequences
$d_{\infty}\text{-convergence}\left\{\begin{aligned} &\implies\text{weak convergence}\\ &\implies d_{1}\text{-convergence}\implies+\text{-convergence}\implies\text{coffecientwise}\end{aligned}\right.$
@@ 5. Proof of the main results
(2) Assume that $\mathfrak{a}_{\bullet,0}\neq\mathfrak{a}_{\bullet,1}$. We first assume that there exists $f\in R$ and $\lambda\in\mathbb{R}_{>0}$ such that $f\in\mathfrak{a}_{\lambda,0}\backslash\mathfrak{a}_{\lambda,1}$ and $f\notin\mathfrak{a}_{\lambda^{\prime},0}$ for $\lambda^{\prime}>\lambda$. Take $t\in(0,1)$ and $\mu,\nu$ with $(1-t)\mu+t\nu=\lambda$. We claim that $f\notin\mathfrak{a}_{\lambda,t}$. Indeed, if $\mu\leq\lambda$, then $\nu\geq\lambda$, so $f\notin\mathfrak{a}_{\nu,1}$ since $f\notin\mathfrak{a}_{\lambda,1}$. If $\mu>\lambda$, then by the choice of $f$ and $\lambda$ we have $f\notin\mathfrak{a}_{\mu,0}$. So in either case, $f\notin\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$. This shows that $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$. If $\mathfrak{a}_{\bullet,1}\nsubseteq\mathfrak{a}_{\bullet,0}$, then we can repeat the above construction to show that $\mathfrak{a}_{\bullet,t}\neq\mathfrak{a}_{\bullet,0}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
Otherwise, $\mathfrak{a}_{\bullet,1}\subsetneq\mathfrak{a}_{\bullet,0}$. We claim that there exists $\lambda\in\mathbb{R}_{>0}$ and $\epsilon>0$ such that $\mathfrak{a}_{\lambda+\epsilon,0}\supsetneq\mathfrak{a}_{\lambda,1}$ and $\mathfrak{a}_{\lambda-\epsilon,1}\supsetneq\mathfrak{a}_{\lambda,1}$. Indeed if all the jumping numbers of $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ coincide then the claim is clear. Otherwise, choose a jumping number $a$ of $\mathfrak{a}_{\bullet,1}$ with $\mathfrak{a}_{a,1}\subsetneq\mathfrak{a}_{a,0}$ which is not a jumping number of $\mathfrak{a}_{\bullet,0}$. Then there exists $a^{\prime}>a$ such that $\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}$. Let $\lambda:=(a+a^{\prime})/2$ and $\epsilon:=\lambda-a$, then $\mathfrak{a}_{\lambda-\epsilon,1}=\lambda_{a,1}\supsetneq\mathfrak{a}_{\lambda,1}$ since $a$ is a jumping number and $\mathfrak{a}_{\lambda+\epsilon,1}=\mathfrak{a}_{a^{\prime},1}=\mathfrak{a}_{a,1}\supsetneq\mathfrak{a}_{a,0}\supset\mathfrak{a}_{\lambda,0}$. Thus we showed $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. Now it follows from Proposition 3.6 that $(\mathrm{Fil}^{s},d_{1})$ is convex.
@@
Note that by 3.3, $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$ is a rooftop metric space ([Xia23, Definition 5.2]. It is easy to verify that a decreasing (resp. increasing) Cauchy sequence $\mathfrak{a}_{\bullet,k}$ in $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ satisfies $\cap_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ (resp. $\cup_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$). Therefore by Proposition 3.15, Lemma 3.25 and [Xia23, Proposition 5.4], $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is complete. ∎
@@
(2) It follows from Definition-Lemma 2.2, Definition-Lemma 2.22 and Proposition 2.24 that $(\mathrm{Fil}^{s},\subset,\cap,\vee_{s})$ is a distributive lattice. By Proposition 2.23, the saturation from $\mathrm{Fil}$ to $\mathrm{Fil}^{s}$ is a joint morphism, which is clearly surjective. By Example 2.25, the map $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\to\mathrm{Fil}^{s}$, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is injective, and is a join morphism as the composition of two join morphisms. ∎
@@ 5.1. Continuity properties of log canonical thresholds
In this section, we assume $(R,\mathfrak{m})$ is a klt singularity over a field $\mathbbm{k}$ of characteristic 00.
@@
$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k}).$ (5.1)
@@
$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}).$ (5.2)
@@
$\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})>c+\epsilon.$ (5.3)
@@
(1) follows immediately from Proposition 5.1.
To see (2), assume that $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology. By [Blu18, Theorem B.1], there exists $v\in\mathrm{Val}^{<+\infty}_{X,x}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet})=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}$. By definition, we have $\lim_{k}v(\mathfrak{a}_{\bullet,k})=v(\mathfrak{a}_{\bullet})$, hence
@@
(3) follows from Proposition 4.6. The proof is finished. ∎

### v2

@@ 4. Supnorm metrics on the space of filtrations
In this section, we study the basic properties of $d_{\infty}$, prove the local Lipschitz continuity of log canonical thresholds, and compare several different topologies on the filtration spaces.
@@ 4.1. Supnorm metrics and homogenization
Throughout this subsection, we will use the language of norms. Recall that $\mathcal{N}$ is the set of linearly bounded $\mathfrak{m}$-norms on $R$.
@@
Fix a norm $\rho\in\mathcal{N}$. We define a function $d_{\infty,\rho}:\mathcal{N}\times\mathcal{N}\to\mathbb{R}_{\geq 0}$ by
@@
It is not hard to see that the $d_{\infty}$-topology is independent of the choice of the reference norm.
@@
By definition, $d_{\infty,\rho}$ is symmetric and non-negative for any $\rho$. Given $\chi,\chi^{\prime},\chi^{\prime\prime}\in\mathcal{N}$, since $|\chi(f)-\chi^{\prime\prime}(f)|\leq|\chi(f)-\chi^{\prime}(f)|+|\chi^{\prime}(f)-\chi^{\prime\prime}(f)|$, the triangle inequality holds. This proves that $d_{\infty,\rho}$ is a pseudometric.
The second claim follows from the fact that given $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\gg 0$ such that $M^{-1}\rho\leq\rho^{\prime}\leq M\rho$. ∎
On $\mathcal{N}^{h}$, we do not need to take $\limsup$ in the definition of $d_{\infty}$.
@@
Denote $d:=d_{\infty}(\widehat{\chi},\widehat{\chi}^{\prime})$. By Lemma 4.3, we may assume that $d=\sup\frac{\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)}{\chi_{0}(f)}$. For any $\epsilon>0$, there exists $f\in\mathfrak{m}$ such that
@@
The following lemma shows that $(\mathcal{N}^{h},d_{\infty})$ is complete.
@@
4.2. Local Lipschitz continuity of log canonical thresholds
In this subsection we prove that the log canonical threshold is locally Lipschitz continuous with respect to the $d_{\infty}$-topology.
@@ 4.2. Local Lipschitz continuity of log canonical thresholds
[…]
@@
(4.1) $\frac{v(f)}{\chi^{\prime}(f)}=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi(f)-\chi^{\prime}(f)}{\chi^{\prime}(f)}\right)$
@@ 4.3. Comparison of different topologies on $\mathrm{Fil}^{s}$
Now we would like to compare the topologies on the space $\mathrm{Fil}^{s}$.  [^3] Since the product topology is in general not first countable and hence not obviously sequential, we avoid using the term finer for them, but only compare convergence for sequences. First we show that a $d_{1}$-converging sequence converges in the ++-topology.
@@
Let $\mathfrak{a}_{\bullet}$$\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ for $k\in\mathbb{Z}_{>0}$. If $\lim_{k}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet})=0$, then $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}$ in the ++-topology.
@@
We claim that if $\mathfrak{a}_{\bullet,k}\subset\mathfrak{b}_{\bullet,k}$ satisfies $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{b}_{\bullet,k})=0$, then for any $v\in\mathrm{Val}^{+}_{X,x}$, we have $\lim_{k\to\infty}(v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{b}_{\bullet,k}))=0$.
@@
Now it remains to prove the claim, which follows essentially from the same argument as [BLQ24, Proposition 3.12]. We include a sketch for the reader’s convenience. Assume to the contrary that there exist $\epsilon>0$ and $v\in\mathrm{Val}^{+}_{X,x}$ such that $v(\mathfrak{a}_{\bullet,k})>v(\mathfrak{b}_{\bullet,k})+\epsilon$ for any $k$. Then for $k\in\mathbb{Z}_{>0}$, we may choose $\ell_{k}\in\mathbb{Z}_{>0}$ and $f_{k}\in\mathfrak{b}_{\ell_{k},k}$ such that
@@
(4.2) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\leq M\cdot d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}).$
@@
To summarize, we have the following diagram for the convergence of sequences
$d_{\infty}\text{-convergence}\left\{\begin{aligned} &\implies\text{weak convergence}\\ &\implies d_{1}\text{-convergence}\implies+\text{-convergence}\implies\text{coefficientwise convergence}\end{aligned}\right$.
@@ 5. Proof of the main results
(2) Since $\mathfrak{a}_{\bullet,0}\neq\mathfrak{a}_{\bullet,1}$, we may assume that $\mathfrak{a}_{\bullet,0}\not\subseteq\mathfrak{a}_{\bullet,1}$. Then there exist $f\in R$ and $\lambda\in\mathbb{R}_{>0}$ such that $f\in\mathfrak{a}_{\lambda,0}\backslash\mathfrak{a}_{\lambda,1}$, and $f\notin\mathfrak{a}_{\lambda^{\prime},0}$ for $\lambda^{\prime}>\lambda$. Take $t\in(0,1)$ and $\mu,\nu$ with $(1-t)\mu+t\nu=\lambda$. We claim that $f\notin\mathfrak{a}_{\lambda,t}$. Indeed, if $\mu\leq\lambda$, then $\nu\geq\lambda$, so $f\notin\mathfrak{a}_{\nu,1}$ since $f\notin\mathfrak{a}_{\lambda,1}$. If $\mu>\lambda$, then by the choice of $f$ and $\lambda$ we have $f\notin\mathfrak{a}_{\mu,0}$. In either case, $f\notin\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$. Thus $f\notin\mathfrak{a}_{\lambda,t}$ and hence $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,0}$.
If $\mathfrak{a}_{\bullet,1}\nsubseteq\mathfrak{a}_{\bullet,0}$, then we can repeat the above construction to show that $\mathfrak{a}_{\bullet,t}\neq\mathfrak{a}_{\bullet,1}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ for any $t\in(0,1)$.
Otherwise, $\mathfrak{a}_{\bullet,1}\subsetneq\mathfrak{a}_{\bullet,0}$. We claim that there exist $\lambda\in\mathbb{R}_{>0}$ and $\epsilon>0$ such that $\mathfrak{a}_{\lambda+\epsilon,0}\supsetneq\mathfrak{a}_{\lambda,1}$ and $\mathfrak{a}_{\lambda-\epsilon,1}\supsetneq\mathfrak{a}_{\lambda,1}$. Indeed, if all the jumping numbers of $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ coincide then the claim is clear. Otherwise, choose a jumping number $a$ of $\mathfrak{a}_{\bullet,1}$ with $\mathfrak{a}_{a,1}\subsetneq\mathfrak{a}_{a,0}$ which is not a jumping number of $\mathfrak{a}_{\bullet,0}$. Then there exists $a^{\prime}>a$ such that $\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}$. Let $\lambda:=(a+a^{\prime})/2$ and $\epsilon:=\lambda-a$, then
$\mathfrak{a}_{\lambda-\epsilon,1}=\mathfrak{a}_{a,1}\supsetneq\mathfrak{a}_{\lambda,1}$
since $a$ is a jumping number of $\mathfrak{a}_{\bullet,1}$, and
$\mathfrak{a}_{\lambda+\epsilon,0}=\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}\supsetneq\mathfrak{a}_{a,1}\supset\mathfrak{a}_{\lambda,1}$
by construction. So we get
$\mathfrak{a}_{\lambda,1}\subsetneq\mathfrak{a}_{\lambda+\epsilon,0}\cap\mathfrak{a}_{\lambda-\epsilon,1}\subset\sum_{\mu+\nu=2\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}=\mathfrak{a}_{\lambda,1/2},$
and hence $\mathfrak{a}_{\bullet,1/2}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. Now it follows from Proposition 3.6 that $(\mathrm{Fil}^{s},d_{1})$ is convex.
Otherwise, $\mathfrak{a}_{\lambda,0}\subsetneq\mathfrak{a}_{\lambda,1}$ for any $\lambda\in\mathbb{R}_{>0}$.
@@
Note that by Lemma 3.3, $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ is a rooftop metric space ([Xia23, Definition 5.2]. It is easy to verify that a decreasing (resp. increasing) Cauchy sequence $\mathfrak{a}_{\bullet,k}$ in $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ satisfies $\cap_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ (resp. $\cup_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$). Therefore by Proposition 3.15, Lemma 3.25 and [Xia23, Proposition 5.4], $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ is complete. ∎
@@
(2) It follows from Definition-Lemma 2.2, Definition-Lemma 2.22 and Proposition 2.24 that $(\mathrm{Fil}^{s},\subset,\cap,\vee_{s})$ is a distributive lattice. By Proposition 2.23, the saturation from $\mathrm{Fil}$ to $\mathrm{Fil}^{s}$ is a join morphism, which is clearly surjective. By Example 2.25, the map $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\to\mathrm{Fil}^{s}$, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is injective, and is a join morphism as the composition of two join morphisms. ∎
@@ 5.1. Continuity properties of log canonical thresholds
In this section, we assume $(R,\mathfrak{m})$ is a klt singularity over a field $\mathbbm{k}$ of characteristic 0.
@@
(5.1) $\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k}).$
@@
(5.2) $\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}).$
@@
(5.3) $\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})>c+\epsilon.$
@@
Assertion (1) follows immediately from Proposition 5.1.
To see assertion (2), assume that $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the ++-topology. By [Blu18, Theorem B.1], there exists $v\in\mathrm{Val}^{<+\infty}_{X,x}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet})=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}$. By definition, we have $\lim_{k}v(\mathfrak{a}_{\bullet,k})=v(\mathfrak{a}_{\bullet})$, hence
@@
Assertion (3) follows from Proposition 4.6. The proof is finished. ∎